A rectangle is 4 times as long as it is wide. The perimeter of a rectangle is 50cm. what are the dimensions of the rectangle?
Answer by macston(5194)

.
W=width; L=length=4W
.
P=2(L+W)
50cm=2(L+W)
25cm=L+W
25cm=4W+W
25cm=5W
5cm=W
.
L=4W=4(5cm)=20cm
.
ANSWER: Length is 20cm and width is 5cm.
.
CHECK:
P=2(L+W)
50cm=2(20cm+5cm)
50cm=2(25cm)
50cm=50cm
.